OpenStudy (anonymous):

Tom wants to hang a mirror in his room. The mirror and frame must have an area of 14 square feet. The mirror is 3 feet wide and 4 feet long. Which quadratic equation can be used to determine the thickness of the frame, x?

OpenStudy (anonymous):

A) 4x2 + 14x - 2 = 0 B) 3x2 + 10x - 14 = 0 C) 2x2 + 12x - 14 = 0 D) x2 + 7x - 14 = 0